# Cosmological Singularity and Power-Law Solutions in Modified Gravity

### v1

@@ Abstract
A bouncing Universe avoids the big-bang singularity. Using the time-like and null Raychaudhhuri equations, we explore whether the bounce near the big-bang, within a broad spectrum of modified theories of gravity, allows for cosmologically relevant power-law solutions under reasonable physical conditions. Our study shows that certain modified theories of gravity, such as Stelle gravity, do not demonstrate singularity resolution under any reasonable conditions, while others including $f(R)$ gravity and Brans-Dicke theory can demonstrate singularity resolution under suitable conditions. For these theories, we show that the accelerating solution is slightly favoured over ekypyrosis.
@@ I Introduction
To study the existence of singularities in general and in particular at the big-bang, we use the Raychaudhuri equation, which as we know, governs the evolution of ‘expansion of geodesics’, and is the primary tool behind the celebrated Hawking-Penrose singularity theorems [3, 4, 5, 6, 7]. The equation predicts the existence of conjugate points, both in the finite past and the finite future. A congruence of geodesics meet at these points, making the tangent vectors ambiguous. This in turn implies the incompleteness of geodesics via a set of straightforward arguments, which is generally accepted as the necessary and sufficient condition for a spacetime to be singular[8]. The only assumption that one needs to make is that the energy-momentum under question satisfies reasonable energy conditions. These can of course be broken under certain circumstances, for example when one has a positive cosmological constant (which has negative pressure) or when certain modified theories of gravity are considered instead of GR.
Our paper is organized as follows. In the next section, we do a quick review of the power-law solutions in cosmology. In section III we remind the reader about the generic conditions of the bounce. In Section IV, we review the Raychaudhuri equation and its prediction of conjugate points and singularities under very general circumstances.
In Sections V, VI and VII, we study bounces in the context of Stelle-gravity, $f(R)$ and Brans-Dicke theory, arguably some of the best studied and well-motivated modifications of GR. We collect our results, summarize and discuss future directions in Section VIII.
@@ II Cosmology and Power-law Solutions
Our objective is to examine whether various modified gravity models result in cosmological power-law solutions, when avoiding the singularity. These solutions serve as cosmological attractors [9]. For example, an exponential potential yields a scale factor with a power-law behavior. Our investigation covers both late-time accelerating solutions and ekpyrotic power-law cosmologies for the early universe. For power-law solutions, the scale factor is as follows:
@@
There exist various types of power-law solutions. The power-law solution leads to accelerated expansion when $\alpha>1$. However, it becomes ekpyrotic when $\alpha<1/3$. This paper specifically concentrates on these two types of power-law solutions. Other values of $\alpha$ correspond to generic power-law solutions, which are of no direct interest to our current work.
@@
III The bounce
In this section, we examine the various conditions that arise if one demands a bouncing cosmological solution, i.e., one which does not start at a singularity. As we know, the expansion of the Universe in GR is governed by the Friedmann equations
$H^{2}\equiv\left(\frac{\dot{a}(t)}{a(t)}\right)^{2}=\frac{8\pi G}{3}\rho$ (2)
and
$\frac{\ddot{a}}{a}=-\frac{4\pi G}{3}\left(\rho+3p\right)$ (3)
where $\rho$ and $p$ are the density and the pressure of the matter content of the Universe respectively. It may be noted that Eq.(3) is also the Raychaudhuri equation in disguise, from which one can obtain the above with the identification of the expansion $\theta\rightarrow 3\dot{a}/a$ [13]. More on the Raychaudhuri equation in the following sections. In the early Universe, most of the matter is in the form of radiation with energy density
$\rho\propto a^{-4}$ (4)
for which the scale-factor according to (2) evolves as
$a(t)\propto t^{\frac{1}{2}}.$ (5)
This implies that
$\rho\propto t^{-2}.$ (6)
Furthermore, the equation of state of radiation, namely $p=\frac{1}{3}\rho$, suggests that the pressure is also time-dependent as follows
$p\propto\frac{1}{3}t^{-2}.$ (7)
As can be seen from Eqs.(5,6.7) above, at time $t=0$ the energy density (6) and the pressure of radiation (7) become infinite while the scale-factor tends to zero, which is known as the big bang singularity. In the standard model of cosmology the big bang is the beginning of time, and the initial state of the universe.
Even though we employed the radiation phase to illustrate the emergence of a singularity, it can be demonstrated that any model more complex than pure de Sitter in GR will inevitably encounter a singularity. For instance, any accelerating solution described by a power-law function with $a\sim t^{\alpha}$, where $\alpha>1$, will not be able to evade this singularity. Another model to capture the accelerating expansion of the early universe is the ekpyrotic model[1], which is characterised by a power-law solution of the form
$a(t)\propto\lvert t\rvert^{\alpha}$ (8)
where $\alpha<\frac{1}{3}$ is a constant. Equation (8) takes negative values of time, ($t<0$) thus allowing us to extrapolate back in time before the big bang. The singularity theorems do not permit such an extrapolation in time, in fact one might have to violate some energy conditions in order for that to occur. The Universe is time-like geodesically incomplete to the past of the big bang. In the case when one has a bounce, the time-like geodesics of the theory can be extended beyond their focal point. When that is so, the big bang is located at the focal point and the infinite density and pressure of the radiation are a manifestation of the presence of a coordinate singularity. If we find coordinates that makes the scale-factor non-zero at time $t=0$, we might be able to extrapolate in the negative time direction without violating the energy conditions. This is the case we are interested in.
IV Raychaudhuri equation and the focusing Theorem
In this section, we study the bounce in the language of the Raychaudhuri equation, which succinctly displays the existence of caustics or conjugate points for a congruence of timelike or null geodesics for most reasonable spacetimes and corresponding stress-energy tensors, which is essential for the singularity theorems to hold. Consequently, conditions which may prevent the formation of such caustics is a useful tool in studying singularity resolution.
Geodesic congruences with timelike and null tangent vector fields are characterized by their expansion $\theta=\nabla_{\mu}u^{\mu}$, $\hat{\theta}=\nabla_{\nu}n^{\nu}$, respectively[14], which satisfy the Raychaudhuri equations
@@ IV Raychaudhuri equation and the focusing Theorem
Next, we will consider three modified gravity models: Stelle gravity, $f(R)$ gravity, and the Brans-Dicke theory. We will investigate whether we can obtain interesting power-law solutions that satisfy either the timelike or null energy condition while potentially avoiding the big-bang singularity.
V Stelle Gravity
@@
V.1 Time-like Raychaudhuri Equation:
@@
It can be shown (see [20]) that the (00) component of the tensor $H_{\mu\nu}$ in FLRW spacetime is given by
@@
where $\lambda=(3\beta-\alpha)/\kappa^{2}>0$, $\alpha$ and $\beta$ are constants of the theory and $H=\frac{\dot{a}}{a}$ is the Hubble parameter. The expansion scalar $\theta$ in FLRW background for the time-like vector $u^{\mu}=(-1,0,0,0)$ can be readily computed
$\theta=3\ \frac{\dot{a}}{a}=\frac{3\alpha}{t},$ (29)
where we used the power-law solution for the scale factor. Thus (26) becomes
@@
V.2 The Null Raychaudhuri Equation
The null Raychaudhuri equation on the other hand takes the form {dmath} d^θdλ=-^θ22 -κ^2ρ_0(w+1)t^-3α(w+1)-2α-12λt^-2(2+α)-3λt^-4(1+α)+t^-2(1+2α). The expansion scalar for null geodesics $(n^{\mu}=\frac{1}{a},\frac{1}{a^{2}},0,0)$ in FLRW spacetime for power-law solution for the scale factor can be readily obtained
$\frac{\hat{\theta}^{2}}{2}=\frac{\alpha^{2}}{2a_{0}^{2}}t^{-2(\alpha+1)}.$ (31)
Plugging-in the above in the null Raychaudhuri equation we get
{dmath}d^θdλ
=-κ^2ρ_0(w+1)t^-3α(w+1)-2α -α22a02t^-2(α+1)-12λt^-2(2+α)-3λt^-4(1+α)+t^-2(1+2α). Since $\lambda>0$, there is only one positive term on the right-hand side of (V.2). However, at very early times, this positive term will not dominate the fourth term, which has the highest negative power of $t$, thereby leaving us with no chance of violating the convergence theorem. In principle, Stelle gravity does not violate singularity theorems, except in cases where $\lambda<0$, which introduces a tachyon into the theory, clearly undesirable [21].
VI f(R) Gravity
@@ V.2 The Null Raychaudhuri Equation
In the following subsection, we will employ 35 for a relatively generic choice of $f(R)$ and examine the power-law solutions that could potentially avoid the singularity.
VI.1 The $R^{n}$ Model
In this paper we will consider an $f(R)$ theory consisting of the usual scalar curvature plus a perturbative correction of higher power in the curvature as $f(R)=R[1+(\ell_{Pl}^{2}R)^{n-1)}]$, where the Planck length is controlling the strength of the corrections, and $n$ is a positive integer. Since we are interested in the large curvature regime, we will consider the case where the second term in $f(R)$ dominates so that $f(R)\sim l^{2(n-1)}R^{n}$. This is the celebrated Starobinsky [23, 24, 25, 26, 27] model. Substituting this $f(R)$ into (35) results in
$3nH^{2}R^{n-1}=\frac{\kappa^{2}\rho}{\ell^{2(n-1)}_{Pl}}+\frac{R^{n}}{2}\left(n-1\right)-3Hn\left(n-1\right)R^{n-2}\dot{R}.$ (36)
This equation will give us a condition relating the power $n$ of the Ricci scalar and the power of the scale factor. Similar to the previous model, we will assume $a(t)\propto t^{\alpha}$. Applying this expression in the Ricci scalar for the FLRW spacetime, which is given by
$R=6\left(\frac{a\ddot{a}+\dot{a}^{2}}{a^{2}}\right)$ (37)
we get the following power-law behaviour
$R\propto t^{-2}.$ (38)
Using this and $\rho=\rho_{0}a^{-3\left(w+1\right)}$ in equation (36) we get the following expression
$t^{-2n}\left(3n\left(2-n\right)-\frac{1}{2}\left(n-1\right)\right)=\frac{\kappa^{2}\rho_{0}}{\ell^{2(n-1)}_{Pl}}t^{-2\epsilon\alpha},$ (39)
where $\epsilon$ is defined to be $\epsilon=\frac{3}{2}\left(w+1\right)$ [28]. This immediately leads to the following condition
$\alpha=\frac{n}{\epsilon}.$ (40)
In the following sections, we will examine the time-like and null Raychaudhuri equations.
VI.2 Time-like Raychaudhuri Equation
The time-like Raychaudhuri equation in $f(R)$ gravity [29] given by
$\frac{d\theta}{d\tau}=-\frac{\theta^{2}}{3}-\frac{\kappa^{2}T_{\mu\nu}u^{\mu}u^{\nu}}{2n\ell_{Pl}^{2n-2}R^{n-1}}+\frac{3}{n}\left(\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}\right)\,$ (41)
can be reduced to {dmath} dθdτ=-3α^2t^-2-κ2ρ0t-3α(w+1)+2n-22n (2α2-α)n-1Pl2n-2
+3n(2α^2-α)t^-2. By using (40) we notice that the time dependence of the second term on the right hand side of the above equation matches with the other terms. Therefore, to violate the convergence condition the sum of the coefficients of $t$ in (VI.2) must be positive:
$\frac{3\alpha(2\alpha-1)}{n}-3\alpha^{2}-\frac{\kappa^{2}\rho_{0}}{2n\ell_{Pl}^{2n-2}\left(2\alpha^{2}-\alpha\right)^{n-1}}>0.$ (42)
The last term in (42) is suppressed by the Planck length and can be ignored. Thus the condition simplifies as
@@ VI.2 Time-like Raychaudhuri Equation
VI.3 The Null Raychaudhuri Equation
The null Raychaudhuri equation in $f(R)$ gravity takes the following form [29],
@@ VI.3 The Null Raychaudhuri Equation
The expression above can be expressed as {dmath} d^θdλ =-12α^2t^-2α-2+(2-2n)(1-2n)nt^-2α-2κ2ρ0(1+w)2nℓPl2n-2(2α2-α)n-1t^-3α(w+1)-2α+2n-2. Just as in the time-like case, the time dependence on the right-hand side of the above equation is identical for all terms. Ignoring the Planck-suppressed term, the condition for violating the focusing theorem simplifies to:
@@
VII Brans-Dicke theory
@@
when we compare the first term on the right hand side of (VII) to the source term in the usual Einstein field equations of GR. The relation between $\phi$ and G in (48) indicates that G varies with time. The coupling constant $\omega$ is a tunable parameter that can be adjusted to fit the data at hand.
VII.1 Dynamics from the Friedmann Equations
@@
The (00) component of the Einstein-Brans-Dicke equation (VII) in a spatially flat FLRW spacetime sourced by a perfect fluid takes the form:
@@
VII.2 Time-like Raychaudhuri Equation
@@
In figure 4 we display the LHS of the inequality (70) for a fixed value of $\omega=-3.5/2$, which is consistent with the other condition 64) and for various values of $\alpha=1.2,2,3$ which are larger than 1. We see that there is an interval of values of $\gamma$ for which the expression in (70 )is positive, hence satisfies the inequality.
@@
Figure 9: Displays the left hand side of the inequality Eq.(74) for various values of $\omega<-3/2$ (consistent with 64).
@@
As in the accelerating case, the largest possible value of $\gamma$ is 2. Now we will plot (Fig. 9) the LHS of the inequality 74 for $\omega<-3/2$ to see if there is any $\gamma$ that satisfies (74).
@@
VII.3 The Null Raychaudhuri Equation
@@
VIII Conclusions
Table 1: Table showing convergence ($\times$) and divergence (check) for the various models and the time-like and null Raychudhuri equations.

### v2

@@ Abstract
A bouncing Universe avoids the big-bang singularity. Using the time-like and null Raychaudhuri equations, we explore whether the bounce near the big-bang, within a broad spectrum of modified theories of gravity, allows for cosmologically relevant power-law solutions under reasonable physical conditions. Our study shows that certain modified theories of gravity, such as Stelle gravity, do not demonstrate singularity resolution under any reasonable conditions, while others including $f(R)$ gravity and Brans-Dicke theory can demonstrate singularity resolution under suitable conditions. For these theories, we show that the accelerating solution is slightly favoured over ekypyrosis.
@@ I Introduction
To study the existence of singularities in general and in particular at the big-bang, we use the Raychaudhuri equation, which as we know, governs the evolution of ‘expansion of geodesics’, and is the primary tool behind the celebrated Hawking-Penrose singularity theorems [3, 4, 5, 6, 7]. The equation predicts the existence of conjugate points, both in the finite past and the finite future. A congruence of geodesics meet at these points, making the tangent vectors ambiguous. This in turn implies the incompleteness of geodesics via a set of straightforward arguments, which is generally accepted as the necessary and sufficient condition for a spacetime to be singular [8]. The only assumption that one needs to make is that the energy-momentum under question satisfies reasonable energy conditions. These can of course be broken under certain circumstances, for example when one has a positive cosmological constant (which has negative pressure) or when certain modified theories of gravity are considered instead of GR [nojiri2011unified, nojiri2017modified].
Our paper is organized as follows. In the next section, we do a quick review of the power-law solutions in cosmology. In Section III, we review the Raychaudhuri equation and its prediction of conjugate points and singularities under very general circumstances.
In Sections IV, V and VI, we study the corrections to the cosmological backgrounds and bounces in the context of Stelle-gravity, $f(R)$ and Brans-Dicke theory, arguably some of the best studied and well-motivated modifications of GR. We collect our results, summarize and discuss future directions in Section VII.
@@ II Cosmology and Power-law Solutions
Our objective is to examine whether various modified gravity models result in cosmological power-law solutions, when avoiding the singularity. These solutions serve as cosmological attractors [9]. For example, an exponential potential yields a scale factor with a power-law behavior. Our investigation covers both late-time accelerating solutions and ekpyrotic solutions for the early universe. For power-law solutions, the scale factor is as follows:
@@
We get accelerated expansion when $\alpha>1$ and ekpyrosis when $\alpha<1/3$. This paper specifically concentrates on these two types of solutions. Other values of $\alpha$ correspond to generic power-law solutions, which are of no direct interest to our current work.
@@
III Raychaudhuri equation and the focusing Theorem
In this section, we study the bounce in the language of the Raychaudhuri equation, which succinctly displays the existence of caustics or conjugate points for a congruence of time-like or null geodesics for most reasonable spacetimes and corresponding stress-energy tensors, which is essential for the singularity theorems to hold. Consequently, conditions which may prevent the formation of such caustics is a useful tool in studying singularity resolution. In particular, one is looking for additional terms in the RHS of the Raychaudhuri equation, which are positive, as these are the ones which can potentially prevent the formation of caustics, and thereby render the singularity theorems inapplicable.
Geodesic congruences with time-like and null tangent vector fields are characterized by their expansion $\theta=\nabla_{\mu}u^{\mu}$, $\hat{\theta}=\nabla_{\nu}n^{\nu}$, respectively[14], which satisfy the Raychaudhuri equations
@@ III Raychaudhuri equation and the focusing Theorem
Next, we will consider the application of three modified gravity models: Stelle gravity, $f(R)$ gravity, and the Brans-Dicke theory cosmological spacetimes. We will investigate whether we can obtain interesting power-law solutions that satisfy either the timelike or null energy condition while potentially avoiding the big-bang singularity.
IV Stelle Gravity
@@
IV.1 Time-like Raychaudhuri Equation:
@@
It can be shown (see [20]) that the (00) component of the tensor $H_{\mu\nu}$ is given by
@@
where $\lambda=(3\beta-\alpha)/\kappa^{2}>0$, $\alpha$ and $\beta$ are constants of the theory and $H=\frac{\dot{a}}{a}$ is the Hubble parameter. Including these results, the equation (19) becomes
@@
IV.2 The Null Raychaudhuri Equation
The null Raychaudhuri equation on the other hand takes the form {dmath} d^θdλ=-κ^2ρ_0(w+1)t^-3α(w+1)-2α -α22a02t^-2(α+1)-12λt^-2(2+α)-3λt^-4(1+α)+t^-2(1+2α). Since $\lambda>0$, there is only one positive term on the right-hand side of (IV.2). However, at very early times, this positive term will not dominate the fourth term, which has the highest negative power of $t$, thereby leaving us with no chance of violating the convergence theorem. In principle, Stelle gravity does not violate singularity theorems, except in cases where $\lambda<0$, which introduces a tachyon into the theory, and that is clearly undesirable [21].
V f(R) Gravity
@@ IV.2 The Null Raychaudhuri Equation
In the following subsection, we will employ (26) for a relatively generic choice of $f(R)$ and examine the power-law solutions that could potentially avoid the singularity.
V.1 The $R^{n}$ Model
In this paper we will consider an $f(R)$ theory consisting of the usual scalar curvature plus a perturbative correction of higher power in the curvature as $f(R)=R[1+(\ell_{Pl}^{2}R)^{n-1)}]$,  [^1] Note the presence of the $l_{Pl}^{2}$ term to ensure the correct dimensions.The Planck length $\ell_{Pl}$ is controlling the strength of the corrections, and $n$ is a positive integer. Since we are interested in the large curvature regime $R\gg 1/\ell^{2}_{Pl}\implies R\ell^{2}_{Pl}\gg 1$, we will consider the case where the second term in $f(R)$ dominates so that $f(R)\simeq R^{n}$. In this paper, this is the $f(R)$ gravity model [23, 24, 25, 26, 27] that we will investigate for singularity avoidance. In the following sections, we will examine the time-like and the null Raychaudhuri equations.
V.2 Time-like Raychaudhuri Equation
The time-like Raychaudhuri equation in $f(R)=R^{n}$ gravity [29] is given by
$\frac{d\theta}{d\tau}=-\frac{\theta^{2}}{3}-\frac{\kappa^{2}T_{\mu\nu}u^{\mu}u^{\nu}}{2n\ell_{Pl}^{2n-2}R^{n-1}}+\frac{3}{n}\left(\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}\right).$ (27)
For Einstein gravity ($n=1$), the second term in the above equation will read
$-\frac{\kappa^{2}T_{\mu\nu}u^{\mu}u^{\nu}}{2n\ell_{Pl}^{2n-2}R^{n-1}}=-\frac{\kappa^{2}}{2}T_{\mu\nu}u^{\mu}u^{\nu},$ (28)
which is not supressed by the large curvature term, unlike the $R^{n}$ case, where the stress energy term becomes subleading due to the presence of the $R^{n-1}$ term in the denominator.
The remaining sum (without the second term) of (27) in Einstein’s gravity reduces to
$-\frac{\theta^{2}}{3}+\frac{3}{n}\left(\frac{\ddot{a}}{a}+3\frac{\dot{a}^{2}}{a^{2}}\right)=3\frac{\ddot{a}}{a}$ (29)
[…]
Now, to violate the convergence condition the sum of the coefficients of $t$ in (V.2) must be positive:
$\frac{3\alpha(2\alpha-1)}{n}-3\alpha^{2}>0.$ (30)
Thus the condition simplifies as
@@ V.2 Time-like Raychaudhuri Equation
V.3 The Null Raychaudhuri Equation
The null Raychaudhuri equation in $f(R)=R^{n}$ gravity takes the following form [29],
@@ V.3 The Null Raychaudhuri Equation
We will see that, just like the time like case, extending the model from Einstein gravity to $R^{n}$ model makes the stress-energy term, subleading, and will introduce the necessary positive term. To see this explicitly, consider $n=1$, then the stress energy term simplifies as follows
$-\frac{\kappa^{2}T_{\mu\nu}n^{\mu}n^{\nu}}{2n\ell_{Pl}^{2\left(n-1\right)}R^{n-1}}\simeq-\frac{\kappa^{2}}{2}T_{\mu\nu}n^{\mu}n^{\nu},$ (33)
which is clearly not suppressed by higher curvature term. On top of that the necessary positive term
$+\frac{1}{R^{n-1}}\left(n^{t}\right)^{2}\partial^{2}_{t}R^{n-1}=0$ (34)
will not exist. The expression (32) above can thus be expressed as {dmath} d^θdλ =-12α^2t^-2α-2+(2-2n)(1-2n)nt^-2α-2. Just as in the time-like case, the time dependence on the right-hand side of the above equation is identical for all terms. Ignoring the Planck-suppressed term, the condition for violating the focusing theorem simplifies to:
@@
VI Brans-Dicke theory
@@
when we compare the first term on the right hand side of (VI) to the source term in the usual Einstein field equations of GR. The relation between $\phi$ and G in (38) indicates that G varies with time. The coupling constant $\omega$ is a tunable parameter that can be adjusted to fit the data at hand.
VI.1 Dynamics from the Friedmann Equations
@@
The (00) component of the Einstein-Brans-Dicke equation (VI) for perfect fluid takes the form:
@@
VI.2 Time-like Raychaudhuri Equation
@@
In figure 4 we display the LHS of the inequality (60) for a fixed value of $\omega=-3.5/2$, which is consistent with the other condition 54) and for various values of $\alpha=1.2,2,3$ which are larger than 1. We see that there is an interval of values of $\gamma$ for which the expression in (60 )is positive, hence satisfies the inequality.
@@
Figure 9: Displays the left hand side of the inequality Eq.(64) for various values of $\omega<-3/2$ (consistent with 54).
@@
As in the accelerating case, the largest possible value of $\gamma$ is 2. Now we will plot (Fig. 9) the LHS of the inequality 64 for $\omega<-3/2$ to see if there is any $\gamma$ that satisfies (64).
@@
VI.3 The Null Raychaudhuri Equation
@@
VII Conclusions
Table 1: Table showing convergence ($\times$) and divergence (check) for the various models and the time-like and null Raychaudhuri equations.